Write a mechanism for the nitration of methyl benzoate (major product only) Include formation of the electrophile from the reaction of nitric acid with sulfuric acid. Only one resonance structure is needed for the intermediate in the EAS portion of the mechanism
Answer #1
Concepts and reason

The concept used to solve this question is to write the mechanism for the nitration of methyl benzoate.

Fundamentals

Nitration of benzene/ substituted benzenes is an electrophilic substitution reaction. The attacking species is an electrophile (NO2+)\left( {{\rm{NO}}_2^ + } \right) . The nitration reaction occurs in the presence of nitration mixture.

Nitration mixture is a mixture of nitric acid and sulfuric acid (HNO3+H2SO4)\left( {{\rm{HN}}{{\rm{O}}_3} + {{\rm{H}}_2}{\rm{S}}{{\rm{O}}_4}} \right) .

In substituted benzenes, the substitution present on the ring determines the position of the attack of the electrophile.

When the substituent is electron-donating species, the electrophilic substitution occurs at ortho/para positions.

When the substituent is electron-withdrawing species, the electrophilic substitution occurs at meta position.

The mechanism is as follows:

H
OH +
6=
=ô
-
HON-0

The mechanism for the attack of electrophile on methyl benzoate is as follows:

CH3
methyl benzoate
NO2

The mechanism for the resonance of carbocation intermediate is as follows:

CH,
--
--المہ
NO2
NO2
NO2

The mechanism for the formation of product is as follows:

CH,
quo
0
-H+
NO2
NO2

Ans:

The mechanism for the nitration of methyl benzoate is as follows:

CH3
-CH
NO2
methyl benzoate
NO2
-CH,
-CH
ph-oh-oh
NO2
NO,
NO2
H+
1
NO2
NO2